# Learning thin deformable object manipulation with a multi-sensory integrated soft hand

### v1

@@ I Introduction
Thin, deformable objects are in many daily activities. For robots to be widely used in human living environments, they must be able to manipulate these objects. Despite notable progress in robotic manipulation [1, 2, 3, 4, 5, 6] and the deployment of numerous commercial robotic platforms [7, 8, 9, 10] with precision surpassing that of humans, even the most advanced robots fall short of human performance when manipulating thin, deformable objects. We suggest that advancements in thin deformable object manipulation hinge on achieving “imprecise dexterity” — the ability to perform dexterous manipulation through robust and adaptive behaviors despite imprecisions in control, movement, and perception.
Singulation (isolating a single layer or item from a group or stack) and grasping thin, deformable objects are common daily tasks, such as turning book pages (Fig. 4 and Fig. 5). These tasks present intricate dynamics due to the thin structures, stacking tendencies, and high-dimensional configuration spaces of the objects, posing significant challenges in manipulation control and perception. Conventional robots, which rely on precise control, struggle to meet multiple objectives: adapting to surface deformation and slipping, maintaining stable contact, avoiding object damage, and ensuring successful singulation and grasping [11]. Visual sensors often fail to capture essential physical properties like friction and stiffness and can be occluded by fingers or objects, necessitating heavy reliance on touch and proprioception for feedback on object interaction. While humans effortlessly perform such tasks with hands that integrate passive compliance, touch, and proprioception [12], it remains an open challenge in robotics.
@@
The key component is our soft, underactuated two-finger hand equipped with tactile and F/T sensors, offering passive compliance, touch, and muscle proprioception similar to a human hand. Different thin, deformable objects exhibit various behaviors under external forces; for instance, fabric denting easily while pages resist pressure but yield to shear forces. Our gripper’s passive compliance allows its fingers to follow the object’s deformation, maintaining continuous contact and limiting potentially damaging forces exerted on both objects and the robot [11]. Such adaptability enables gentle, force-limited interactions, resulting in successful grasps without requiring perfect robot precision [22, 23]. The underactuated design reduces control parameters and simplifies the manipulation, while the integrated sensors provide the foundation for measuring the environmental information necessary for the task.
To drive our soft hand, we learned a direct mapping from perception to robot control using model-free RL in the real world. However, model-free RL alone was insufficient for learning a robust singulation and grasping policy. Therefore, we incorporated and validated several additional components to achieve the demonstrated capabilities.
@@
The first is a slip module. We found the reward sparsity in these tasks makes robot impossible to discover effective policies from scratch, and acquiring task-relevant perception remains a problem. To address these, we employ an active exploratory motion called slip, where robot fingers slide over the object’s surface. Slip motion is guided by a neural network that receives inputs from a wrist-mounted camera and is trained on data reflecting human intuition (e.g., turning book pages from the upper right). Such position-specific slips provide the robot with a proper initial pose, reducing the need for exploration. Meanwhile, the deformation brought by the interaction reveals unobservable physical properties like flexural rigidity and friction, which are captured by the tactile and F/T sensors.
@@
Another component is a hierarchical, dual-loop learning structure. In essence, the dual-loop structure lies in the structured and efficient exploration and exploitation of the action space. The necessary action range after slip motion varies due to differences in material surface properties and movement errors caused by depth camera noise. The outer loop selects between coarse and fine action spaces based on a required range of movement, while the inner loop outputs concrete robot actions. By selecting the appropriate action space based on current perception and learned experience, unnecessary action exploration is reduced.
The presented study achieves a significant step toward human-level sensorimotor intelligence. Our key contribution is an approach for singulation and grasping thin, deformable objects by controlling a soft, underactuated hand that integrates passive compliance and multimodal perception. The resulting approach enables adaptive and dexterous manipulation without requiring high precision or prior knowledge, as demonstrated by experiments. The singulating and grasping of thin, deformable objects have long been difficult problems in robotics. Our methodology provides a general framework that addresses this challenge and paves the way for downstream robotic manipulation tasks.
@@
Our prior work [14] presented a preliminary study restricted to manipulating paper-like objects. Here, we propose new methods (Section III) to expand the manipulation scope to a wide range of thin, deformable objects. This includes a new problem description (Section III-A), integration of tactile sensors (Section III-B), a new slip module (Section III-C) for obtaining sensory information, a double-loop learning structure (Section III-D), and a multisensory encoder (Section III-H) to enhance learning efficacy and performance. We conduct extensive experiments (Section IV) demonstrating the effectiveness of our method on challenging tasks beyond those covered in our previous work, while also elaborating on and evaluating the role of passive compliance (Section IV-C and Section IV-D) and multisensory integration (Section IV-F). A comprehensive component-wise analysis is provided (Section IV-E), along with a discussion on limitations and future opportunities (Section IV-G), offering a broader perspective on the applicability and impact of our approach.
@@ II Related Work
Conventional approaches to deformable object manipulation have revolved around accurately modeling the object’s dynamics and subsequently applying model predictive control or trajectory optimization [24, 25, 26, 27, 28]. These methods heavily depend on the precise information about the object and any uncertainties encoded in the dynamics model being accurate [29, 30, 31]. Some research attempts to model a single thin, deformable object as points or a line based on empirically tuned values, which are often unreliable [32]. Overall, these systems are limited to scenarios that can be precisely modeled and lack robustness against the diverse uncertainties encountered in the singulation and grasping of thin, deformable objects.
@@
For manipulating thin, deformable objects, model-free reinforcement learning (RL) methods are promising alternatives [33, 34, 35, 36, 37] to bypass traditional dynamics modeling. Model-free RL features by learning to control robots directly from raw sensory data through continuous trial and error, introducing a paradigm shift in many robotic areas [38, 39, 40, 41, 42, 43]. The manipulation of deformable objects using RL has made significant advances in environments with abundant simulation data [41, 42]. However, applying model-free RL to learn singulate and grasp thin, deformable objects remains challenging. These objects’ sensitivity to external forces requires the common rigid robot hands to precisely control the contact between fingers and the object, as well as manage the amount and timing of the applied force [11]. This creates a high-dimensional control space that is difficult to learn with RL and poses risks of damaging both the robot and the object during real-world exploration. The challenge is exacerbated by the inability of simulators, commonly used to accelerate RL training, to replicate the dynamics of thin, deformable objects accurately [43]. Successful learning to manipulate these objects may require simplifying manipulation while ensuring adaptability to varying object properties without relying on high-precision control.
@@
Soft robotic hands have recently gained attention for simplifying manipulation while adapting to varying object properties without the need for high-precision control [44, 14, 11]. Studies show that soft hands offer passive stability and make manipulations easier to achieve[22, 23]. Meanwhile, the inherent compliance of soft robotic hands handles uncertainties during object-fingertip engagement gracefully and limits the potentially damaging forces exerted on thin, deformable objects [11].
@@
Figure 2: Overview of the presented approach. (A) We train manipulation policies in the real world using RL. During training, the robot proactively interacts with the object to obtain multisensory observations via a slip module. The outer loop selects an action space based on observations, and the inner loop explores within it to output concrete actions for the robot to execute. After training, we deploy the learned policy zero-shot on robots for various tasks. (B) Slip module. The object mask, derived from the segment anything model, is fed into a fully convolutional network to infer pixel-wise affordance maps. The chosen pixel is decoded into slip motion.
@@ III Methods
First, the robot obtains multisensory observations through the slip module (Fig. 2B). We use an RGB-D camera to obtain visual observations, allowing for the inference of the slip point and its direction by the slip module. Subsequently, the robot touches and slips with objects and monitors through tactile sensors and force-torque sensors. Thus, the obtained multisensory observations provide insights into unobservable physical information of the environment state.
@@
Then, the outer loop selects action space from two granularity levels, coarse and fine, based on the multisensory observations. The coarse action space offers broader movements with sparse discretization, while the fine action space covers a smaller range but benefits from dense discretization. The outer loop utilizes the multisensory encoder to integrate the observations from different senses into a latent vector representation, which then guides the action space selection.
@@
Finally, the inner loop explores and learns within the selected action space. Employing the same multisensory encoder, the inner loop maps the observations to the concrete action that is executed by the robot. During training, the rewards obtained by the inner loop are shared with the outer loop, achieving mutual performance improvement.
III-A Problem formulation
We formulate the problem of singulation and grasping thin, deformable objects as a Markov Decision Process (MDP), with a state space S, action space $A$, a scalar reward function $R$, and a transition probability $P(s_{t+1}|s_{t},a_{t})$. At time step $t$, an agent uses a policy $\pi$ to select an action $a_{t}$ and receives rewards $r_{t}=R(s_{t},a_{t})$.
Specifically, in our inner-outer loop learning process, the policy function is composed of two components, the outer loop policy $\pi^{out}$ and the inner loop policy $\pi^{in}$. Given the current state, the agent selects a range of action $a^{out}$ using $\pi^{out}$. Then, given $a^{out}$, the agent predicts concrete action $a^{in}$ for the robot to execute using $\pi^{in}$. The objective is to learn an optimal joint policy $(\pi^{out*},\pi^{in*})$ that maximizes the expected discounted reward $r_{t}$, which can be written as
$(\pi^{out*},\pi^{in*})=\arg\max_{\begin{subarray}{c}\pi^{out}\pi^{in}\end{subarray}}E\left[\sum_{t=0}^{\infty}\lambda^{t}R\left(s_{t},\pi^{in}\left(s_{t}\mid a^{out}_{t}\right)\right)\right],$
where ${E}[\cdot]$ is the expectation after acting according to the optimal policy in the environment, and $\lambda$ is the discount factor. The reward function $R$ is binary, assigning a value of 1 for task success and 0 otherwise.
III-B Hardware setup
@@
Fig. 3A provides an overview of the hardware setup used for policy training. The setup includes two Universal Robot 5 robot arms, each equipped with an Intel Realsense L515 camera, an ATI gamma Force Torque sensor at the wrist joint, and a two-finger soft robotic hand.
@@
The soft robotic hand is based on the design of the fast-PneuNet [53], a simple pneumatic actuator widely used in the research community. The whole hand is made of thermoplastic polyurethane (TPU) using a consumer-grade 3D printer. The index and thumb fingertips of each hand are equipped with a vision-based tactile sensor [54]. Furthermore, the F/T sensor and the soft robotic hand are fixed together, moving and rotating in unison. Section S4 provides further details about tactile sensors and the soft gripper.
@@
We maneuver the robot using position control within the local/tool coordinate system $H_{\alpha\beta\gamma}$, which is established at the index fingertip, as shown in Fig. 3C. The $\alpha$ axis is oriented along the line that connects the two fingertips and lies within the longitudinal plane $\Lambda$ formed by two fingers; the $\gamma$ axis is orthogonal to the $\alpha$ axis and remains within plane $\Lambda$; the $\beta$ axis is perpendicular to the plane formed by $\alpha$ and $\gamma$. Each command issued to the robot is interpreted as movements relative to its current pose. Compared with directly controlling joint angles of the robot arm, this allows us to simplify the dimensions of the action in reinforcement learning and reduce the difficulty of policy learning.
III-C Slip module
The slip module serves two primary functions: obtaining task-related sensory information and providing an initial robot pose to minimize unnecessary exploration. To obtain rich multisensory observations, we employ a slip motion to proactively interact with the object (Fig. 2B). This process begins with the robot capturing aligned color and depth images from the top of the object. The color image is processed through the Segment Anything Model (SAM) [55] to obtain object masks. These masks are rotated at various angles and fed to the slip network to return a set of affordance maps. Each map contains values corresponding to the slip motion primitive, which are parameterized by the pixel’s location and its rotation angle within the map. The pixel with the highest affordance value is selected; its position identifies the slip point, and the corresponding rotation angle of the affordance map determines the slip direction.
We next perform the slip motion, where the robot’s hand rotates so that the $\alpha$ axis, i.e., the line connecting the two fingers, aligns with the slip direction. The index finger then approaches the object’s surface and moves to the predicted slip point. Subsequently, positive air pressure is applied to the soft gripper, enabling the fingers to contact and interact with the object. In this process, we adjust the pose of the end effector to ensure that the plane composed of the $\alpha$ and $\beta$ axes is parallel to the workspace. During the training phase, the workspace’s inclination is maintained at 0 degrees. Meanwhile, the position of the slip point in the camera coordinate system is obtained via the camera’s depth channel. We beforehand established the relation between the camera and the robot’s fingertip coordinate system through hand-eye calibration.
After a slip, we record 3D arrays of deformation flows from the tactile sensors as touch observation, along with readings from the F/T sensor as the muscle proprioception, including forces $(f{\alpha},f{\beta},f{\gamma})$ and torques $(m{\alpha},m{\beta},m{\gamma})$. The three main axes of the F/T sensor are aligned with the three corresponding axes of the $H_{\alpha\beta\gamma}$. This alignment makes the force and torque data measured by the sensor more intuitive. Moreover, we crop the area surrounding the slip point on the depth image to serve as visual observation.
Through the robot’s purposive movements prior to manipulation, observations are made that maximize the information uptake on the relevant properties of the target object. These observations provide insights into the elaborate interplay between static and dynamic sensory cues, such as texture, softness, and friction.
The slip network is pre-trained using synthetic data before the robot learns in the real world. Humans annotated slip points on a few images featuring a book, a shirt, and a spring pancake. We employed data augmentation techniques, including rotation, translation, and scaling, on the object masks of these images to simulate different object sizes and positions. More details of the slip network architecture are in Section S6.
@@ III-C Slip module
III-D Manipulation policy training via inner-outer loops
During training, within the inner-outer loop policy learning framework, we aim to find an action space selection policy $\pi^{out}$ using the outer loop, and an optimal manipulation policy $\pi^{in}$ under the selected action space using the inner loop. Both loops use the same multisensory observations, and the information about action space selection is included as part of the observations of the inner loop. We utilize Soft Actor-Critic (SAC) [56] for learning the policies, where the reward obtained by the inner loop is shared with the outer loop during each iteration.
III-E Observation and action
The state for both inner and outer loop policy is defined as $s_{t}=(o_{t}^{vis},o_{t}^{ind},o_{t}^{thu},o_{t}^{pro},o_{t}^{aux})$, where $o_{t}^{vis}$ refers to the visual observation from the camera, $(o_{t}^{ind},o_{t}^{thu})$ refers to the touch observation from the index and thumb fingertips, respectively; $o_{t}^{prop}$ refers to the muscle proprioception from the F/T sensor, and $o_{t}^{aux}$ refers to the auxiliary state. The auxiliary state includes the action space selection and a flag representing whether the current stage belongs to the inner or outer loop. In the outer loop, the flag is assigned a placeholder value of zero to maintain uniformity in the observation format, thereby facilitating the reuse of the multisensory encoder by the inner loop.

### v2

@@ I Introduction
Thin, deformable objects are in many daily activities. For robots to be widely used in human living environments, they must be able to manipulate these objects. Despite notable progress in robotic manipulation [1, 2, 3] and the deployment of numerous commercial robotic platforms, such as Covariant’s AI-driven picking solutions [4] and Boston Dynamics’ Stretch warehouse robot [5], even the most advanced robots fall short of human performance when manipulating thin, deformable objects. We suggest that advancements in thin deformable object manipulation hinge on achieving “imprecise dexterity” — the ability to perform dexterous manipulation through robust and adaptive behaviors despite imprecision in control, movement, and perception.
Singulation and grasping thin, deformable objects are common daily tasks, such as turning book pages (Fig. 1). These tasks present intricate dynamics due to the thin structures, stacking tendencies, and high-dimensional configuration spaces of the objects, posing significant challenges in manipulation control and perception. Conventional robots, which rely on precise control, struggle to meet multiple objectives: adapting to surface deformation and slipping, maintaining stable contact, avoiding object damage, and ensuring successful singulation and grasping [6]. Visual sensors often fail to capture essential physical properties like friction and stiffness, and can be occluded by fingers or objects, necessitating heavy reliance on touch and proprioception for feedback on object interaction. While humans effortlessly perform such tasks with hands that integrate passive compliance, touch, and proprioception [7], it remains an open challenge in robotics.
@@
One of the key components is our soft, underactuated two-finger hand. Different thin deformable objects exhibit various behaviors under external forces; for instance, fabric dents easily while pages resist pressure but yield to shear forces. Our gripper’s passive compliance allows its fingers to follow the object’s deformation, maintaining continuous contact and limiting potentially damaging forces exerted on both objects and the robot [6]. Such adaptability enables gentle, force-limited interactions, resulting in successful grasps without requiring perfect robot precision [17, 18]. Compared with the gripper in Flipbot [9], we maintained its general design but augmented its sensory capabilities. Our hand integrates tactile sensors at each fingertip alongside the original F/T sensors, offering passive compliance, touch, and muscle proprioception similar to a human hand. The addition of tactile sensing enables the robot to perceive local contact features closely linked to material properties such as texture and hardness. In contrast, Flipbot’s reliance solely on global force information provided by F/T sensors limits its performance and generalization potential, restricting it to the paper-like object.
Like Flipbot [9], we learned a direct mapping from perception to robot control using model-free RL in the real world to drive our soft hand. Model-based methods, while powerful in structured domains, struggle in deformable object manipulation due to the inherent uncertainty and complexity of thin deformable objects. Their high-dimensional configuration spaces, coupled with partially observable dynamics and large variations in material properties (e.g., bending stiffness, friction), pose significant challenges for analytical modeling. In contrast, model-free RL learn manipulation policies end-to-end from trial-and-error exploration, without requiring explicit modeling of object dynamics. However, model-free RL alone was insufficient for learning the robust policy as presented in our video. Therefore, we incorporated and validated several additional components to achieve the demonstrated capabilities.
@@
The first is a slip module. We found that the reward sparsity in these tasks makes robot impossible to discover effective policies from scratch, and acquiring task-relevant perception remains a problem. To address these, we employ an active exploratory motion called slip, where robot fingers slide over the object’s surface. In Flipbot [9], the slip motion is hardcoded and fixed for specific scenarios, making it impossible to generalize to different object sizes and positions. Now, slip motion is guided by a neural network that receives inputs from a wrist-mounted camera and is trained on data reflecting human intuition (e.g., turning book pages from the upper right). Such position-specific slips not only provide the robot with a proper initial pose, reducing the search space for policy learning, but also generate rich, dynamic interactions that expose latent physical properties like flexural rigidity and friction.
@@
Another component is a hierarchical, dual-loop learning structure, designed to improve learning efficiency—an essential factor for real-world reinforcement learning. Flipbot [9] introduced a reset mechanism and an automatic reward signal based on the page number, significantly reducing human intervention and enabling policy learning within four hours. Building on this framework, we extend the training setup to include fabric materials. This increased variety necessitates greater exploration and longer training durations. To address this, the dual-loop structure accelerates learning through the structured and efficient exploration and exploitation of the action space. The necessary action range after slip motion varies due to differences in material surface properties and movement errors caused by depth camera noise. The outer loop selects between coarse and fine action spaces based on a required range of movement, while the inner loop outputs concrete robot actions. By selecting the appropriate action space based on current perception and learned experience, unnecessary action exploration is reduced.
The presented study achieves a significant leap over the existing work in terms of graspable object variety and performance, marking a meaningful step toward human-level sensorimotor intelligence. Notably, our training only uses printer paper and winter fabric on flat surfaces. Yet when deployed in environments, the robot successfully handles various deformable objects with different physical properties (coated paper, hotel towel, summer suit fabric), hybrid materials (fabrics separated by the introduction page or spring pancakes separated by baking paper), and gravitational effects. The singulating and grasping of thin, deformable objects have long been difficult problems in robotics. Our results suggest that the extraordinary complexity of manipulating thin, deformable objects can be tamed with the integration of embodied multi-sensory integration and passive compliance.
The main contributions of this article are as follows:
  1. 1.
    We propose a novel approach for singulation and grasping thin, deformable objects by controlling a soft, underactuated hand that integrates passive compliance and multi-modal perception. To the best of our knowledge, this is the first work to successfully demonstrate the singulation and grasping across such a wide range of thin deformable objects under challenging real-world conditions.
  2. 2.
    We introduce a slip module that actively collects rich multisensory information and provides a favorable initial pose prior to manipulation. This enhances task awareness and learning performance.
  3. 3.
    We present a dual-loop learning structure that adaptively selects action space granularity based on sensory input. The mechanism improves sample efficiency and accelerates policy convergence.
  4. 4.
    We validate our approach through extensive experiments, including ablation studies, quantitative evaluations of passive compliance, and analysis of the underlying mechanisms enabling successful manipulation.
@@
This journal paper evolves upon the conference paper presented in [9]. Section III introduces the problem formulation (Section III-B), a slip module (Section III-C) for obtaining sensory information, and a double-loop learning structure. We conduct extensive experiments (Section IV) demonstrating the effectiveness of our method, while also evaluating the role of passive compliance (Section IV-C and Section IV-D) and multisensory integration (Section IV-F). A comprehensive component-wise analysis is provided (Section IV-E), along with a discussion on limitations and future opportunities (Section IV-G).
@@ II Related Work
Deformable object manipulation encompasses a range of strategies. One line of research involves model-based methods, which often revolve around accurately modeling the object’s dynamics and subsequently applying model predictive control or trajectory optimization [19, 20, 21]. These methods heavily depend on the precise information about the object and any uncertainties encoded in the dynamics model being accurate [22, 23, 24]. In the context of thin, deformable objects, model-based approaches face significant challenges due to multiple sources of uncertainty. Accurately estimating physical properties such as bending stiffness and surface friction is inherently difficult, as these parameters vary even among seemingly identical objects due to differences in texture, wear, and manufacturing variability. Moreover, such objects exhibit high-dimensional and nonlinear configuration spaces, where small action perturbations result in drastically different deformations, rendering reliable forward prediction intractable. These uncertainties are further exacerbated by sensor occlusion, depth perception errors, and complex finger-object interactions, which are difficult to incorporate into analytical models.
Another line of research that has gained substantial attention for deformable object manipulation is model-free approaches. Notably, Navarro-Alarcon et al. [25] proposed a model-free visual servoing framework to control the deformation of elastic objects using closed-loop feedback, bypassing the need for explicit object modeling. Concurrently, learning-based methods have prominence in deformable object manipulation [26, 27], which is largely attributed to the capability to infer action from sensory input, bypassing the complexities of direct physical modeling [28, 29, 30, 31].
@@
Within this domain, model-free reinforcement learning (RL) methods [32, 33] have shown great potential for manipulating thin, deformable objects. Model-free RL features by learning to control robots directly from raw sensory data through continuous trial and error, introducing a paradigm shift in many robotic areas, including whole-body control for door opening [34], planetary exploration using legged robots [35], and autonomous flight of micro aerial swarms [36]. The manipulation of deformable objects using RL has made significant advances in environments with abundant simulation data [37, 38]. Moreover, sim-to-real transfer techniques have further extended these capabilities to real-world applications [39]. However, applying model-free RL to learn thin, deformable object manipulation remains challenging. These objects’ sensitivity to external forces requires the common rigid robot hands to precisely control the contact between fingers and the object, as well as manage the amount and timing of the applied force [6]. This creates a high-dimensional control space that is difficult to learn with RL and poses risks of damaging both the robot and the object during real-world exploration. The challenge is exacerbated by the inability of simulators, commonly used to accelerate RL training, to replicate the dynamics of thin, deformable objects accurately. Blanco-Mulero et al. [40] systematically analyzed this gap, revealing that state-of-the-art simulators struggle to accurately replicate the nuanced physical behaviors of cloth. Successful learning to manipulate these objects may require simplifying manipulation while ensuring adaptability to varying object properties without relying on high-precision control.
@@
Soft robotic hands [41, 42, 43] have recently gained attention for simplifying manipulation while adapting to varying object properties without the need for high-precision control [44, 6]. Studies show that soft hands offer passive stability and make manipulations easier to achieve [17, 18]. Meanwhile, the inherent compliance of soft robotic hands handles uncertainties during object-fingertip engagement gracefully and limits the potentially damaging forces exerted on thin, deformable objects [6].
@@
Figure 2: Overview of the presented approach. (A) We train manipulation policies in the real world using RL. During training, the robot proactively interacts with the object to obtain multisensory observations via a slip module. The outer loop selects an action space based on observations, and the inner loop explores within it to output concrete actions for the robot to execute. After training, we deploy the learned policy zero-shot on robots for various tasks. (B) Slip module. The object mask, derived from the segment anything model, is fed into a fully convolutional network to infer pixel-wise affordance maps. The chosen pixel is decoded into slip motion, with its 3D coordinates derived from the depth image.
@@ III Methods
First, the robot obtains multisensory observations through the slip module (Fig. 2B). The slip module is pre-trained using data reflecting human preferences. We use an RGB-D camera to obtain visual observations, allowing for the inference of the slip point and its direction by the slip module. Subsequently, the robot touches and slips with objects and monitors through tactile sensors and force-torque sensors. Thus, the obtained multisensory observations provide insights into unobservable physical information of the environment state.
@@
Then, the outer loop selects the action space from two granularity levels, coarse and fine, based on the multisensory observations. The coarse action space offers broader movements with sparse discretization, while the fine action space covers a smaller range but benefits from dense discretization. The outer loop utilizes the multisensory encoder to integrate the observations from different senses into a latent vector representation, which then guides the action space selection.
@@
Finally, the inner loop explores and learns within the selected action space. Employing the same multisensory encoder, the inner loop maps the observations to the robot’s concrete action. During training, the rewards obtained by the inner loop are shared with the outer loop, achieving mutual performance improvement.
III-A Hardware setup
@@
Fig. 3A provides an overview of the hardware setup used for policy learning. The setup includes two Universal Robot 5 robot arms, each equipped with an Intel RealSense L515 camera, an ATI gamma Force Torque sensor at the wrist joint, and a two-finger soft robotic hand.
@@
The soft robotic hand is based on the design of the fast-PneuNet [53], a simple pneumatic actuator widely used in the research community. The whole hand is made of thermoplastic polyurethane (TPU) using a consumer-grade 3D printer. The index and thumb fingertips of each hand are equipped with a vision-based tactile sensor [54]. Furthermore, the F/T sensor and the soft robotic hand are fixed together, moving and rotating in unison. Section S4 provides further details about tactile sensors, and we discuss our gripper design in Section S9.
@@
We maneuver the robot using position control within the local/tool coordinate system $H_{\alpha\beta\gamma}$, which is established at the index fingertip, as shown in Fig. 3C. The $\alpha$ axis is oriented along the line that connects the two fingertips and lies within the longitudinal plane $\Lambda$ formed by two fingers; the $\gamma$ axis is orthogonal to the $\alpha$ axis and remains within plane $\Lambda$; the $\beta$ axis is perpendicular to the plane formed by $\alpha$ and $\gamma$. Each command issued to the robot is interpreted as movements relative to its current pose. Compared with directly controlling the joint angles of the robot arm, this allows us to simplify the dimensions of the action and reduce the difficulty of policy learning.
@@
III-B Problem formulation
We formulate the problem of singulation and grasping thin, deformable objects as a Markov Decision Process (MDP), with a state space $S$, action space $A$, a scalar reward function $R$, and a transition probability $P(s_{t+1}|s_{t},a_{t})$. Given our model-free RL setup, the transition dynamics $P$ are not explicitly modeled but instead learned implicitly through trial-and-error interactions with the environment. At time step $t$, an agent uses a policy $\pi$ to select an action $a_{t}$ and receives rewards $r_{t}=R(s_{t},a_{t})$.
Specifically, we adopt a hierarchical policy learning framework consisting of two interrelated loops: an outer loop policy $\pi^{out}$ and an inner loop policy $\pi^{in}$. Given the current state, the agent selects an action granularity $a^{out}$ using $\pi^{out}$. Then, given $a^{out}$, the agent predicts a concrete action $a^{in}$ for the robot to execute using $\pi^{in}$. The objective is to learn an optimal joint policy $(\pi^{out*},\pi^{in*})$ that maximizes the expected cumulative reward:
$(\pi^{out*},\pi^{in*})=\mathop{\arg\max}_{\pi^{out},\pi^{in}}\;\\ E\bigg{[}\sum_{t=0}^{\infty}\lambda^{t}\,R\big{(}s_{t},\pi^{in}(s_{t}\mid\pi^{out}(s_{t}))\big{)}\bigg{]},$ (1)
where ${E}[\cdot]$ represents the expectation under the joint policy $(\pi^{out*},\pi^{in*})$, and $\lambda$ is the discount factor. The reward function $R$ is binary, assigning a value of 1 for task success and 0 otherwise. We utilize Soft Actor-Critic (SAC) [55] for learning the policies. The definitions of state, action, and reward are as follows.
III-B1 State
The state for both inner and outer loop policy is defined as $s_{t}=(o_{t}^{vis},o_{t}^{ind},o_{t}^{thu},o_{t}^{pro},o_{t}^{aux})$, where $o_{t}^{vis}$ refers to the visual observation from the camera, $(o_{t}^{ind},o_{t}^{thu})$ refers to the touch observation from the index and thumb fingertips, respectively; $o_{t}^{pro}$ refers to the muscle proprioception from the F/T sensor, and $o_{t}^{aux}$ refers to the auxiliary state. The auxiliary state includes the action space selection and a flag representing whether the current stage belongs to the inner or outer loop. In the outer loop, the flag is assigned a placeholder value of zero to maintain uniformity in the observation format, thereby facilitating the reuse of the multisensory encoder by the inner loop.
III-B2 Action